# Embodied Co-Design for Rapidly Evolving Agents: Taxonomy, Frontiers, and Challenges

### v1

@@ Abstract
Brain-body co-evolution enables animals to develop complex behaviors in their environments. Inspired by this biological synergy, embodied co-design (ECD) has emerged as a transformative paradigm for creating intelligent agents—from virtual creatures to physical robots—by jointly optimizing their morphologies and controllers rather than treating control in isolation. This integrated approach facilitates richer environmental interactions and robust task performance. In this survey, we provide a systematic overview of recent advances in ECD. We first formalize the concept of ECD and position it within related fields. We then introduce a hierarchical taxonomy: a lower layer that breaks down agent design into three fundamental components—controlling brain, body morphology, and task environment—and an upper layer that integrates these components into four major ECD frameworks: bi-level, single-level, generative, and open-ended. This taxonomy allows us to synthesize insights from more than one hundred recent studies. We further review notable benchmarks, datasets, and applications in both simulated and real-world scenarios. Finally, we identify significant challenges and offer insights into promising future research directions. A project associated with this survey has been created at https://github.com/Yuxing-Wang-THU/SurveyBrainBody.
@@ I Introduction
Embodied intelligence (EI) seeks to integrate cognitive capabilities into autonomous machines, enabling them to learn from and adapt to complex environments [howard2019evolving]. Situated at the intersection of artificial intelligence and robotics, EI has recently made significant advancements in fusing perception, decision-making, and actuation. This progress has empowered various robots to tackle previously intractable challenges, such as dynamic locomotion [lee2020learning, zhang2024whole], end-to-end navigation [zheng2024towards, schumann2024velma], and dexterous manipulation [chi2023diffusion, huang2023voxposer, wang2024dexcap], to name just a few.
@@
In traditional EI, an agent’s morphology (body) is typically hand-designed and treated as a fixed part of the environment, while only its controller (brain) is optimized for task-oriented behaviors. However, this separation has clear limitations. Since an agent’s performance is intertwined with both its morphology and control, a pre-specified morphology can be suboptimal: it constrains sensing and actuation, thereby narrowing the space of achievable behaviors. In other words, when morphological design prioritizes mechanical optimality without considering how it will be controlled, the burden of achieving coordinated behavior shifts almost entirely to the controller. This imbalance often becomes a bottleneck for task performance and raises a fundamental question: How should we design the optimal form of an embodied agent to effectively perform tasks in its environment?
Figure 1: Differences between control design only and embodied co-design (ECD).
@@
The concept of morphological computation [paul2006morphological, muller2017morphological, pfeifer2009morphological, pfeifer2006body, mertan2024no] posits that a well-designed body can offload or simplify computations typically assigned to the brain, thereby facilitating more efficient behaviors. This perspective introduces a brain-body dynamic: changes in morphology necessitate corresponding adaptations in control to exploit newly acquired capabilities, while environmental feedback can, in turn, guide further refinement of morphology. These synergies have motivated various efforts to improve the current design procedure for embodied agents. As a result, embodied co-design (ECD), an integrated design framework that co-optimizes an agent’s morphology with its control mechanism to enhance task performance, becomes a critical topic in EI.
@@
ECD was originally studied in the Evolutionary Robotics (ER) community [harvey2005evolutionary, silva2016open], where both robotic behaviors and morphologies are concurrently optimized using Evolutionary Algorithms (EAs). However, this topic remains emergent for EI, particularly as agents and their tasks continue to grow in complexity and diversity. In contrast to controller-only optimization, ECD introduces a vast and often combinatorial design space for morphology-control configurations, accompanied by fragile interdependencies. Specifically, evaluating a candidate morphology typically requires re-training the controller from scratch, since even minor morphological changes can invalidate previously learned controllers. Moreover, variations in environmental conditions, design constraints, and task objectives produce heterogeneous feedback signals, further complicating the co-design process, as illustrated in Fig. 1. Consequently, ECD is highly non-convex and computationally expensive, highlighting the increasing necessity for automated ECD techniques over traditional manual engineering.
Figure 2: Co-designed embodied agents in both virtual and real environments. Example images numbered from (a) to (p) are adapted from [sims1994competition, hornby2001evolving, shim2003generating, Veenstra2023, ma2021diffaqua, ha2019reinforcement, luck2020data, zhao2020robogrammar, gupta2021embodied, xu2021accelerated, cheney2013unshackling, zardini2021seeking, medvet2021biodiversity, wang2023curriculum, wang2023softzoo, yue2025toward], respectively. Example images numbered from (q) to (y) are adapted from [nygaard2021environmental, revolve2, Strgar-RSS-24, xie2024map, schaff2022soft, liu_voxcraft_2020, kobayashi2024computational, matthews2023efficient, cheng2024structural], respectively.
@@
Motivation: Recent breakthroughs in domains such as evolutionary algorithms [davis2023subtract, bergonti2024co, ferigo2025totipotent], reinforcement learning (RL) [yuan2022transformact, wang2023curriculum, dongleveraging2024, lu2025bodygen], differentiable simulation [xu2021accelerated, ma2021diffaqua, Strgar-RSS-24], and large models [gao2025vlmgineer, fang2025robomore, song2025laser] have led to a surge of novel ECD approaches, rapidly reshaping the capabilities of co-design systems. However, their assumptions, representations, computational pipelines, evaluation benchmarks, and limitations have not been systematically reviewed, which makes it difficult for researchers to develop a coherent understanding of the ECD landscape. A comprehensive survey is increasingly needed. Our goal is to consolidate these frontier developments, clarify conceptual foundations, highlight emerging trends, and identify open challenges that will shape the next generation of embodied intelligence research.
@@
Taxonomy: To enhance readability for researchers across disciplines, we propose a two-layer hierarchical taxonomy in Section III. The lower layer structures ECD around three core dimensions: controlling brain (Section III-A), body morphology (Section III-B), and task environment (Section III-C). Building upon this foundation, the upper layer categorizes existing ECD methods into four major frameworks: bi-level (Section IV-A), single-level (Section IV-B), generative (Section IV-C), and open-ended (Section IV-D), with further subcategories identified within each. This taxonomy establishes a common vocabulary and provides a conceptual map for embodied co-design, facilitating cross-domain understanding. It also encourages researchers to explore beyond their primary areas. For instance, specialists in RL for control optimization may discover how their methodologies align with a specific ECD framework, in conjunction with previously unconsidered morphological evolution, thereby opening pathways for new interdisciplinary collaborations.
@@
Scope: ECD bridges robotics, artificial intelligence, control theory, and cognitive science. Guided by our proposed taxonomy, we focus on analyzing the design spaces and representation methods of morphology, control, and task environments. More importantly, we investigate how these components are jointly coordinated within different co-design frameworks. Our aim is to provide an organizational perspective that clarifies the conceptual and technical interrelations among various ECD works, rather than exhaustively reviewing any particular algorithmic subfield (e.g., RL for control optimization) or engineering detail (e.g., 3D printing for morphological fabrication).
@@
Related Works: Most prior surveys either provide broad overviews of embodied intelligence [long2025survey, duan2022survey, sun2024comprehensive, liu2025embodied] or focus on specific subtopics, such as embodied learning [liang2025large], manipulation [li2025survey], and navigation [xiong2025sensing]. These works primarily emphasize sensorimotor learning, while ECD has received limited attention. Partial reviews of ECD exist within specific domains, including soft robotics [zhao2024exploring, chen2020design, stroppa2024optimizing], humanoid robots [liu2025embracing, yue2025toward], evolutionary robotics [doncieux2014beyond, prabhu2018survey, alattas2019evolutionary, pandey2022accessible, strgar2024evolution], and artificial life [lai2021virtual, casadei2023artificial, ha2022collective]. However, they lack a structured analytical framework for systematically comparing co-design methodologies and do not incorporate recent advances, benchmarks, and emerging applications. In contrast, our survey provides a comprehensive and up-to-date analysis of ECD, covering a wide range of agent types and offering in-depth discussions on key challenges and future directions.
Contributions: To the best of our knowledge, this paper presents the first comprehensive survey of ECD. The main contributions are summarized below:
@@
    We conceptualize modeling approaches to ECD from three complementary perspectives: bi-level optimization, sequential decision-making, and model-based optimization (Section II). Together, this set of lenses provides a multifaceted view of the ECD landscape.
@@
    We contribute a hierarchical taxonomy (Section III) that organizes co-design spaces and identifies key design choices. The proposed taxonomy aims to deepen researchers’ understanding of the fundamental components involved in embodied co-design.
@@
    We present a systematic technical review (Section IV) of major advances in ECD over the past five years, analyzing representative methodologies and their core principles.
@@
    We survey simulation platforms and real-world testbeds commonly used in ECD research (Section V), highlighting their capabilities, representative demonstrations, and application domains.
@@
    We identify current bottlenecks and existing gaps in ECD, outlining promising research directions (Section VI) that may benefit a broad range of communities and help advance the state of the art.
The remainder of this survey is organized as follows. We begin with the background of ECD in Section II, which includes an introduction to its fundamental elements, mathematical definitions, and related topics. In Section III, we introduce our hierarchical taxonomy. In Section IV, we elaborate on four co-design frameworks based on the proposed taxonomy. We discuss notable benchmarks and contemporary applications of ECD in Section V. Finally, we outline the challenges of modern ECD and provide future perspectives in Section VI.
@@ II-A1 Embodied Agents
An embodied agent, such as a robot, can be characterized by its body morphology and controlling brain. The body morphology $\omega\in\Omega$ defines the agent’s physical interface with the outer surroundings, including its structural topology, material properties, and the spatial arrangement of sensors and actuators. The controlling brain $\pi\in\Pi$ maps observed states to action commands, driving the body to exert forces that interact with the task environment. These forces induce state transitions governed by the physical dynamics $F\in\mathcal{F}$, which, in turn, generate new sensory inputs. Through this sensorimotor loop, the agent learns to produce inherently goal-directed behaviors, with its performance evaluated using a fitness function $R\in\mathcal{R}$.
@@ II-A3 Sequential Decision-Making View
While the bi-level formulation captures the hierarchical structure inherent in ECD, it fails to specify the causal connection between morphology and control. To bridge this gap, we formulate ECD as a single-level Markov Decision Process (MDP) [wang2022deep]. Following the sequential workflow commonly adopted in engineering practice, we divide the overall co-design process into two main stages: the design stage $\mathcal{D}$ and the control stage $\mathcal{C}$. As a result, the ECD problem can be represented as a $9$-tuple $\left\langle\mathcal{S_{D}},\mathcal{S_{C}},\mathcal{A_{D}},\mathcal{A_{C}},\mathcal{F_{D}},\mathcal{F_{C}},\mathcal{R_{D}},\mathcal{R_{C}},\gamma\right\rangle$, where $\gamma\in(0,1]$ denotes the reward discount factor, and the remaining elements correspond to the relevant design spaces. For clarity, explicit representations of constraints—such as morphological or task-specific constraints—are omitted and incorporated implicitly within their respective design spaces.
In this formulation, the overall state space comprises two components: the design state $s_{t}^{d}\in\mathcal{S_{D}}$, which governs morphology creation, and the control state $s_{t}^{c}\in\mathcal{S_{C}}$, which governs interactions with the environment. The corresponding action space includes design actions $a_{t}^{d}\in\mathcal{A_{D}}$ and control actions $a_{t}^{c}\in\mathcal{A_{C}}$. Under an episodic framework, a morphology is initially generated as $\omega=s_{T_{d}}^{d}$ by a design policy $\pi_{d}(a_{t}^{d}|s_{t}^{d})$, which executes a sequence of design actions guided by the design transition dynamics $F_{d}(s^{d}_{t+1}|s^{d}_{t},a^{d}_{t})$ from the initial design state $\omega_{0}=s^{d}_{0}$ to the final design step $T_{d}$. In the subsequent control stage, a policy $\pi_{c}(a_{t}^{c}|s_{t}^{c},\omega)$ conditioned on $\omega$ directs the generated morphology to perform tasks, following the environmental transition dynamics $F_{c}(s^{c}_{t+1}|s^{c}_{t},a^{c}_{t},\omega)$. Based on the above notations, the single-level objective $R$ of ECD is now expressed as:
@@
where $r^{d}_{t}\in\mathcal{R_{D}}$ and $r^{c}_{t}\in\mathcal{R_{C}}$ denote the rewards obtained during the design and control stages, respectively. Leveraging effective gradient approximation techniques in RL [schulman2017proximal, silver2014deterministic, ciosek2020expected], both the design policy and the control policy can be updated iteratively through trial-and-error interactions.
@@ II-A4 Model-Based Optimization View
It is noteworthy that, in the perspectives discussed above, the environmental transition model $F_{c}(s^{c}_{t+1}|s^{c}_{t},a^{c}_{t},\omega)$ is typically treated as a black box. However, as discussed in Section IV-B3, many ECD approaches construct $F_{c}$ as either a differentiable simulator [hu2019chainqueen, du2021diffpd, hu2019difftaichi] or a learned dynamics model [navez2025modeling]. Conceptually, the transition can be written as an abstract function $s^{c}_{t+1}=F_{c}(s^{c}_{t},a^{c}_{t},\omega)$, where $a^{c}_{t}$ denotes a vector of control signals, such as the joint torques applied during that time-step. Thus, given an initially parameterized morphology $\omega$ and a sequence of actions potentially derived from a control policy, the simulator unrolls a trajectory to a terminal control state $s^{c}_{T_{c}}$ and computes a task loss $\mathcal{L}$ that depends on this final state (e.g., distance traveled or final velocity [ma2021diffaqua, xu2021accelerated]). Since the simulator or learned model is differentiable with respect to environmental dynamics, morphological parameters, and control variables, gradients can propagate through the entire trajectory via the chain rule:
@@
This model-based formulation facilitates end-to-end, gradient-based optimization for inverse design in ECD and provides substantial advantages in sample efficiency.

### v2

@@ Abstract
Brain-body co-evolution enables animals to develop complex behaviors in their environments. Inspired by this biological synergy, embodied co-design (ECD) has emerged as a transformative paradigm for creating intelligent agents—from virtual creatures to physical robots—through the joint optimization of morphology and control. This integrated approach facilitates richer environmental interactions and enhances task performance. In this survey, we provide a systematic overview of recent advances in ECD. We first formalize the concept of ECD and position it within related fields. We then introduce a hierarchical taxonomy in which a lower layer breaks down agent design into three fundamental components—controlling brain, body morphology, and task environment—while an upper layer integrates these components into four major ECD frameworks: bi-level, single-level, generative, and open-ended. This taxonomy provides a unified lens through which we distill insights from over one hundred recent studies. We further review notable benchmarks, datasets, and applications in both simulated and real-world scenarios. Finally, we discuss significant challenges and offer insights into promising future research directions. A project associated with this survey has been established at https://github.com/Yuxing-Wang-THU/SurveyBrainBody.
@@ I Introduction
Embodied intelligence (EI) seeks to integrate cognitive capabilities into autonomous machines, enabling them to learn from and adapt to complex environments [howard2019evolving]. Situated at the intersection of artificial intelligence and robotics, EI has achieved significant advancements in fusing perception, decision-making, and actuation, empowering various robots to tackle previously intractable challenges, such as dynamic locomotion [lee2020learning, zhang2024whole], end-to-end navigation [zheng2024towards, schumann2024velma], and dexterous manipulation [chi2023diffusion, huang2023voxposer, wang2024dexcap], to name just a few.
@@
In traditional EI, an agent’s morphology (body) is typically hand-designed and treated as a fixed component of the environment, while only its controller (brain) is optimized for task execution. However, this separation presents clear limitations. A pre-specified morphology can constrain sensing and actuation, thereby narrowing the space of achievable behaviors. In other words, when morphological design prioritizes mechanical optimality without considering how it will be controlled, the burden of achieving coordinated behavior shifts almost entirely to the controller. This imbalance often creates a bottleneck in task performance and raises a fundamental question: How should we design the optimal form of an embodied agent to effectively perform tasks in its environment?
Fig. 1: Differences between control design only and embodied co-design.
@@
The concept of morphological computation [paul2006morphological, muller2017morphological, pfeifer2009morphological, pfeifer2006body, mertan2024no] posits that a well-designed body can simplify computations typically assigned to the brain, thereby facilitating more efficient behaviors. This perspective introduces a brain-body dynamic: changes in morphology necessitate corresponding adaptations in control to exploit newly acquired capabilities, while environmental feedback can, in turn, guide further refinement of morphology. These synergies have motivated various efforts to improve the current agent design procedure. As a result, embodied co-design (ECD), an integrated framework that co-optimizes an agent’s morphology with its control mechanism to enhance task performance, becomes a critical topic in EI.
@@
ECD was initially explored in the Evolutionary Robotics (ER) community [harvey2005evolutionary, silva2016open], where both robot morphology and behavior are jointly optimized using Evolutionary Algorithms (EAs) [harvey2005evolutionary]. However, this topic remains emergent for EI, particularly as agents and their tasks continue to increase in complexity and diversity. Compared to controller-only optimization, ECD introduces a vast, often combinatorial design space for morphology-control configurations, accompanied by fragile interdependencies. Specifically, evaluating a candidate morphology typically requires re-training the controller from scratch, as even minor morphological changes can invalidate previously learned controllers. Variations in environmental conditions, design constraints, and task objectives produce heterogeneous feedback signals, further complicating the co-design process (Fig. 1). Consequently, ECD problems are highly non-convex and computationally expensive, underscoring the growing need for automated methodologies instead of traditional manual engineering.
@@
Motivation: Recent breakthroughs in domains such as evolutionary algorithms [davis2023subtract, bergonti2024co, ferigo2025totipotent], reinforcement learning (RL) [yuan2022transformact, wang2023curriculum, dongleveraging2024, lu2025bodygen], differentiable simulation [xu2021accelerated, ma2021diffaqua, Strgar-RSS-24], and large models [gao2025vlmgineer, fang2025robomore, song2025laser] have led to a surge of novel ECD approaches that rapidly reshape the capabilities of co-design systems. However, their assumptions, representations, computational pipelines, evaluation benchmarks, and limitations have not been systematically reviewed, making it difficult for researchers to develop a coherent understanding of the ECD landscape. Our goal is to clarify the conceptual foundations, consolidate frontier developments, highlight emerging trends, and identify open challenges that will shape the next generation of research in EI.
@@
Taxonomy: To enhance readability for researchers across disciplines, we propose a two-layer hierarchical taxonomy in Section III. The lower layer structures ECD around three core dimensions: controlling brain (Section III-A), body morphology (Section III-B), and task environment (Section III-C). Building upon this foundation, the upper layer categorizes existing ECD methods into four major frameworks: bi-level (Section IV-A), single-level (Section IV-B), generative (Section IV-C), and open-ended (Section IV-D), with further subcategories identified within each. This taxonomy establishes a common vocabulary and provides a conceptual map for ECD, facilitating cross-domain understanding. It also encourages researchers to explore beyond their primary areas. For instance, specialists in RL for control optimization may discover how their methodologies align with a specific ECD framework, in conjunction with previously unconsidered morphological evolution, thereby opening pathways for new collaborations.
@@
Scope: Embodied co-design bridges robotics, artificial intelligence, control theory, and cognitive science. Guided by our proposed taxonomy, we analyze the design spaces and representation methods for morphology, control, and task environments, with a particular emphasis on how these components are coordinated within various co-design frameworks. Rather than exhaustively reviewing any specific algorithmic subfield (e.g., RL for control optimization) or engineering detail (e.g., 3D printing for morphological fabrication), this work provides an organizational perspective that clarifies the conceptual and technical interrelations among different ECD works.
@@
Related Works: Most prior surveys either provide broad overviews of embodied intelligence [long2025survey, duan2022survey, sun2024comprehensive, liu2025embodied] or focus on isolated subfields such as embodied learning [liang2025large], manipulation [li2025survey], and navigation [xiong2025sensing]. These works emphasize sensorimotor integration, while attention to ECD has been largely limited. To our knowledge, partial reviews of ECD exist in specific domains, including soft robotics [zhao2024exploring, chen2020design, stroppa2024optimizing], humanoid robots [liu2025embracing, yue2025toward], evolutionary robotics [doncieux2014beyond, prabhu2018survey, alattas2019evolutionary, pandey2022accessible, strgar2024evolution], and artificial life [lai2021virtual, casadei2023artificial, ha2022collective]. However, these reviews lack a structured taxonomy for systematically comparing ECD methods; they do not incorporate recent advances, benchmarks, and emerging applications.
Features: Unlike previous works, our survey is the first to provide a systematic, method-centric analysis of ECD across various agent types, optimization paradigms, and application domains. Rather than treating morphology as a secondary concern or focusing on isolated subfields, we, for the first time, introduce a two-layer hierarchical taxonomy that explicitly unifies the design process of embodied agents, thus enabling principled comparisons across various ECD methods. We further incorporate recent advances—including differentiable simulation, quality-diversity search, latent and foundation-model-based generation—together with benchmarks, datasets, and real-world testbeds that have been largely absent from related reviews. By combining conceptual unification with practical evaluation criteria and open challenges, this survey offers both a coherent theoretical framework and an up-to-date technical reference for the ECD community.
Contributions: The main contributions of our survey are summarized below:
@@
    We conceptualize ECD from three complementary perspectives: bi-level optimization, sequential decision-making, and model-based optimization (Section II), providing a multifaceted view of the ECD problem.
@@
    We contribute a hierarchical taxonomy (Section III) that organizes co-design spaces and identifies key design choices, offering a unified framework for understanding the fundamental components of ECD.
@@
    We present a systematic technical review (Section IV) of major advances in ECD over the past five years, analyzing representative methods and their core principles.
@@
    We survey simulation platforms, datasets, and real-world testbeds in ECD research (Section V), highlighting their capabilities and representative demonstrations.
@@
    We identify current bottlenecks in ECD and outline promising research directions (Section VI) that may benefit a broad range of communities and help advance the state of the art.
@@ II-A1 Embodied Agents
An embodied agent, such as a robot, can be characterized by its body morphology and controlling brain. The body morphology $\omega\in\Omega$ defines the agent’s physical interface with its environment, encompassing structural topology, material properties, and the spatial configuration of sensors and actuators. The controlling brain $\pi\in\Pi$ maps observed states to action commands, driving the body to exert forces that interact with the task environment. These forces induce state transitions governed by the physical dynamics $F\in\mathcal{F}$, which, in turn, generate new sensory inputs. Through this sensorimotor loop, the agent learns to produce inherently goal-directed behaviors, with its performance evaluated using a fitness function $R\in\mathcal{R}$.
@@ II-A3 Sequential Decision-Making View
While the bi-level formulation captures the hierarchical structure inherent in ECD, it fails to specify the causal connection between morphology and control. To bridge this gap, we formulate ECD as a single-level Markov Decision Process (MDP) [wang2022deep]. Following the sequential workflow commonly adopted in engineering practice, we divide the overall co-design process into two main stages: the design stage $\mathcal{D}$ and the control stage $\mathcal{C}$. As a result, the ECD problem can be represented as a $9$-tuple $\left\langle\mathcal{S_{D}},\mathcal{S_{C}},\mathcal{A_{D}},\mathcal{A_{C}},\mathcal{F_{D}},\mathcal{F_{C}},\mathcal{R_{D}},\mathcal{R_{C}},\gamma\right\rangle$, where $\gamma\in(0,1]$ denotes the reward discount factor, with the remaining elements corresponding to the relevant design spaces. For clarity, explicit representations of constraints—such as morphological or task-specific constraints—are omitted and incorporated implicitly within their respective design spaces.
In this formulation, the overall state space comprises two components: the design state $s_{t}^{d}\in\mathcal{S_{D}}$, which governs morphology creation, and the control state $s_{t}^{c}\in\mathcal{S_{C}}$, which governs interactions with the environment. The corresponding action space includes design actions $a_{t}^{d}\in\mathcal{A_{D}}$ and control actions $a_{t}^{c}\in\mathcal{A_{C}}$. Under an episodic framework, a morphology is initially generated as $\omega=s_{T_{d}}^{d}$ by a design policy $\pi_{d}(a_{t}^{d}|s_{t}^{d})$, which executes a sequence of design actions guided by the design transition dynamics $F_{d}(s^{d}_{t+1}|s^{d}_{t},a^{d}_{t})$ from the initial design state $\omega_{0}=s^{d}_{0}$ to the final design step $T_{d}$. In the subsequent control stage, a policy $\pi_{c}(a_{t}^{c}|s_{t}^{c},\omega)$ conditioned on $\omega$ directs the generated morphology to perform tasks, following the environmental transition dynamics $F_{c}(s^{c}_{t+1}|s^{c}_{t},a^{c}_{t},\omega)$. Based on the aforementioned notations, the single-level objective $R$ of ECD is now expressed as:
@@
where $r^{d}_{t}\in\mathcal{R_{D}}$ and $r^{c}_{t}\in\mathcal{R_{C}}$ denote the rewards obtained during the design and control stages, respectively. By leveraging effective gradient approximation techniques in RL [schulman2017proximal], both the design policy and the control policy can be iteratively updated through trial-and-error interactions.
@@ II-A4 Model-Based Optimization View
It is noteworthy that, in the perspectives discussed above, the environmental transition model $F_{c}(s^{c}_{t+1}|s^{c}_{t},a^{c}_{t},\omega)$ is typically treated as a black box. However, as discussed in Section IV-B3, many ECD approaches construct $F_{c}$ either as a differentiable simulator [hu2019chainqueen, du2021diffpd, hu2019difftaichi] or as a learned dynamics model [navez2025modeling]. Conceptually, the transition can be written as an abstract function $s^{c}_{t+1}=F_{c}(s^{c}_{t},a^{c}_{t},\omega)$, where $a^{c}_{t}$ denotes a vector of control signals, such as the joint torques applied during that time-step. Thus, given an initially parameterized morphology $\omega$ and a sequence of actions potentially derived from a control policy, the simulator unrolls a trajectory to a terminal control state $s^{c}_{T_{c}}$ and computes a task loss $\mathcal{L}$ that depends on this final state (e.g., distance traveled or final velocity [ma2021diffaqua, xu2021accelerated]). Since the simulator or learned model is differentiable with respect to environmental dynamics, morphological parameters, and control variables, gradients can propagate through the entire trajectory via the chain rule:
@@
This model-based formulation facilitates end-to-end, gradient-based optimization for inverse design in ECD, providing substantial advantages in sample efficiency.